OpenStudy (vshiroky):

given the function h described by h(x)= 7/2x+2, find each of the following function values. a. h(0) b. h(-2) c. h(-7) d. h(y)

OpenStudy (mertsj):

Do you know how to substitute?

OpenStudy (vshiroky):

i understood a but not the rest

OpenStudy (mertsj):

How did you get a?

OpenStudy (vshiroky):

I plugged them in, but the other number gives weird fractions

OpenStudy (mertsj):

Weird fractions are just fine.

OpenStudy (vshiroky):

They won't work into my assignment though

OpenStudy (mertsj):

How can they not?

OpenStudy (vshiroky):

I don't know it just wasn't working

OpenStudy (mertsj):

\[h(x)=\frac{7}{2x+2}\] \[h(0)=\frac{7}{2(0)+2}=\frac{7}{2}\]\ \[h(-2)=\frac{7}{2(-2)+2}=\frac{7}{-2}\] \[h(-7)=\frac{7}{2(-7)+2}=\frac{7}{-12}\] \[h(y)=\frac{7}{2(y)+2}\]

OpenStudy (vshiroky):

For b I got -5 and it was right

OpenStudy (mertsj):

\[h(x)=\frac{7}{2}x+2\]

OpenStudy (mertsj):

Is that it?

OpenStudy (vshiroky):

the answer was -5

OpenStudy (vshiroky):

The one I am on now is k(-1)=1/2(-1)+6

OpenStudy (mertsj):

That would be -1/2+6=5 1/2